The volume of a gas is 250 ml at 720 torr and 25°C . If the pressure is held constant and the temperature is raised to 100°C , what will be its new volume?


1
Expert's answer
2021-11-08T08:11:35-0500

V1/T1= V2/T2

250/298=V1/373

V1= 373×0.83893

= 312.92ml